OpenStudy (anonymous):

hello....can someone tell me what happens if the directrix of a parabola is inclined???...

OpenStudy (amistre64):

isnt the directrix perpendicular to the parabola?

OpenStudy (amistre64):

if the directrix becomes inclined, then the parabola is inclined. and we get an xy term in the process

OpenStudy (anonymous):

Hey....thanks for replying.....The definition of a parabola is like its a curve who's points are equidistant from a given point and a given line....

OpenStudy (amistre64):

it can be like that, in fact theres a general conic form that i cant remember that would allow for inclines

OpenStudy (amistre64):

mathically: given a focus(u,v) and a directrix (y=mx+b) then all points that are equidistance can be determined, but its a real pain to work out on here lol

OpenStudy (anonymous):

k man....ill give it a try....and come on.....math isn't a pain...:)

OpenStudy (amistre64):

math aint, but typing is :)

OpenStudy (amistre64):

one method is to simply take a parabola: y = a(x-h)^2+k and run it thru a rotation matrix

OpenStudy (amistre64):

x^ = x cos(t) - y sin(t) y^ = x sin(t) + y cos(t) etc ...

OpenStudy (anonymous):

ok...but my focus is still on the axis....it will get shifted wouldn't it?

OpenStudy (e.mccormick):

The other term you can researc this under is an oblique directrix. That gets some book results.

OpenStudy (anonymous):

hey e.mccormick...thanks...ill check dat out..:)

OpenStudy (amistre64):

or general conic form http://staff.argyll.epsb.ca/jreed/math30p/conics/general.htm

OpenStudy (amistre64):

when the xy term zeros out, the conic is parallel to the x or y axis

OpenStudy (anonymous):

ok...:)